Two supplementary angles are 2x and 4x +108 find x
Answer by cleomenius(959)

2x + 4x + 108 = 180
6x = 72
x = 12.
2(12) = 24 degrees, The smaller angle.
4(12) + 108 = 48 + 108 = 156 degrees, the larger angle.
Cleomenius.